OpenStudy (anonymous):

1)f(x)=2x^(4x) 2)f(x)=4^xlog(small 7)(x) 3)f(x)=[Inx]^4 4)f(x)=In(sqrt(6x+7)(7x-3)) 5)f(x)=5cos(7In(x)) 6(f(x)=x^2(x-7)^5/(x^2+6)^2 f(x)'=

OpenStudy (anonymous):

@mathsmind

OpenStudy (anonymous):

this is homework due tomorrow and i am stuck!

OpenStudy (anonymous):

well they are all simple i can solve them in my head, but let me show u how to do them

OpenStudy (anonymous):

Thank you so much, thats graceful

OpenStudy (anonymous):

solving for Q1 \[f(x)=2x^{4x}\]

OpenStudy (anonymous):

step one: take the natural log of both sides, i will define f(x) as being y ok

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

\[\ln[y=2x^{4x}] \longrightarrow lny= \ln2x^{4x}\]

OpenStudy (anonymous):

clear so far

OpenStudy (anonymous):

\[ lny= 4x(\ln2x)\]

OpenStudy (anonymous):

now we will be using the product rule

OpenStudy (anonymous):

for the right hand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

D(lny)=1/y

OpenStudy (anonymous):

left is 1/y right is 4x(1/2x)+(In2x)(4)

OpenStudy (anonymous):

d(4x(ln2x))= 4ln(2x)=2x^4x(4ln(x)+4)

OpenStudy (anonymous):

=y(2x^4x(4ln(x)+4))

OpenStudy (anonymous):

but we know that y = to the main function

OpenStudy (anonymous):

=2x^(4x)(2x^4x(4ln(x)+4))

OpenStudy (anonymous):

that's the end of q1

OpenStudy (anonymous):

f(x)=4^xlog(small 7)(x)

OpenStudy (anonymous):

u mean to the base 7 right?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

the answer for the first question is \[f(x)= 2x^4(2x^{4x})(4In(x)+4)? \]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

could you write it down by using Equation bottom?

OpenStudy (anonymous):

thanks for your patience.

OpenStudy (anonymous):

\[2x^{4x}(4\ln(x)+4)\]

OpenStudy (anonymous):

i entered it but it says In is not defined

OpenStudy (anonymous):

entered what?

OpenStudy (anonymous):

the answer 2x^(4x)(4In(x)+4)

OpenStudy (anonymous):

yes that is exactly what i did

OpenStudy (anonymous):

undefined because it can't read latex prbably

OpenStudy (anonymous):

probably*

OpenStudy (anonymous):

\[2x^{4x}(4\ln(x)+4)\]

OpenStudy (anonymous):

yes prob, how do i calculate In(x)?

OpenStudy (anonymous):

keep it as it is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I want to know why it wont accpet the answer!

OpenStudy (anonymous):

could you help me with the next one

OpenStudy (anonymous):

if u are submitting ur answer via a text box in a form then write it exactly in the format they ask u to write it in

OpenStudy (anonymous):

look i will solve the problem on a piece of paper and give u the answers ok

OpenStudy (anonymous):

Thank you soooo much for helping out here

OpenStudy (anonymous):

2)f(x)=4^xlog(small 7)(x)

OpenStudy (anonymous):

\[y=4^x \log_{7} (x)\] please confirm

OpenStudy (anonymous):

Yes thats right

OpenStudy (anonymous):

\[= \frac{2\ln(2)4^xln(x)}{\ln(7)}+\frac{4^x}{xln(7)}\]

OpenStudy (anonymous):

i am sorry for rushing this up because there are others waiting for an answer for their problems...

OpenStudy (anonymous):

3)f(x)=[Inx]^4

OpenStudy (anonymous):

\[=\frac{4\ln(x)^3}{x}\]

OpenStudy (anonymous):

its okay thank you im writing the notes what you have just said

OpenStudy (anonymous):

4)f(x)=In(sqrt(6x+7)(7x-3))

OpenStudy (anonymous):

=(1/2)*(6/(7*x-3)-(7*(6*x+7))/(7*x-3)^2)*(7*x-3)/(6*x+7) for Q4)

OpenStudy (anonymous):

f(x)=5cos(7In(x))

OpenStudy (anonymous):

could you use equation bottom for num 4?

OpenStudy (anonymous):

f'(x)= -35*sin(7*ln(x))/x

OpenStudy (anonymous):

for #4 i got everything divided by 6x+7

OpenStudy (anonymous):

6(f(x)=x^2(x-7)^5/(x^2+6)^2

OpenStudy (anonymous):

\[f(x)=\frac{x^2(x-7)^5}{(x^2+6)^2}\]

OpenStudy (anonymous):

please confirm

OpenStudy (anonymous):

Q6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2*x*(x-7)^5/(x^2+6)^2+5*x^2*(x-7)^4/(x^2+6)^2-4*x^3*(x-7)^5/(x^2+6)^3

OpenStudy (anonymous):

man that was a long one hehehehe

OpenStudy (anonymous):

end wow! fin done no more questions for today hehehe

OpenStudy (anonymous):

for this i will give u a medal they are useful derivatives

OpenStudy (anonymous):

thank you so much/ wow it is wrong answer!

OpenStudy (anonymous):

no it is not the wrong answer hehehehe

OpenStudy (anonymous):

show ur work to ur teacher on the piece of paper, and see what she says

OpenStudy (anonymous):

i mean long haha. one more thing, in the end it says x^2+6)^3 it should be power 2?

OpenStudy (anonymous):

= ((x-7)^4 x (3 x^3+14 x^2+42 x-84))/(x^2+6)^3

OpenStudy (anonymous):

i simplified the answer for u

OpenStudy (anonymous):

it doesn't matter because u can solve this in different ways, and simplify both answers i gave u are the same

OpenStudy (anonymous):

i am 100% sure my answers are correct so its up to u

OpenStudy (anonymous):

((x-7)^4 x (3 x^3+14 x^2+42 x-84))/(x^2+6)^3 id it 4x or 4 times?

OpenStudy (anonymous):

there is no 4x its to x-7 to the power 4